Luzadis Company makes furniture using the latest automated technology. The company uses a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The predetermined a cost formula that estimates $1,207000 of total manufacturing overhead for an estimated activity level of During the year, a large quantity of furniture on the market resulted i companys warehouse. The companys cost records revealed the following actual cost and operating data for the year: overhead rate was based on n cutting back production and a buildup of furniture in the Machine-hours Manufacturing overhead cost Inventories at year-end: 55,000 $ 1,169,000 Raw materials Work in process (includes overhead applied of $46,750) Pinished goods (includes overhead applied of $187,000) 11,000 100,000 400,000 1,500,000 Cost of gooda sold (includes overhead applied of $701,250) Required 1. Compute the underapplied or overapplied overhead. 2. Asume that the company closes any underapplied or overaplied overhead to Cost of Goods Sold.Prepare the approplate ds Sold. Prepare the appropriate journal entry 3. Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry 4. How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold? Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4
Answer #1

Requirement #1)

Predetermined Overhead rate =

Total estimated manufacturing OH÷Total estimated machine hrs

=1207000/71000

=$17 per machine hr

Under-applied OH= Actual OH-(Predetermined OH rate × Actual hrs)

=1169000-(17*55000)

=$234,000

Requirement #2)

Cost of goods sold. Dr. 234000

Manufacturing OH. 234,000

(To record Under-applied manufacturing OH)

​​​​​​

Requirement #3)

Calculation of apportionment of Under-applied OH is shown below :

Particulars Overhead applied Ratio Under-applied OH apportioned
Work in process 46750 46750/935000=0.05 234000*0.05=11,700
Finished Goods 187000 187000/935000=0.2 234,000 *0.2=46,800
Cost of goods sold 701250 701250/935000=0.75 234,000 *0.75=175,500.   
Total        935,000 234000

Journal entry :

Work in process 11700

Finished Goods 46800

Cost of goods sold 175500

Manufacturing OH 234,000

(To allocate under applied OH)

Requirement #4

If the underapplied overhead is allocated rather than closed directly to Cost of Goods Sold, the net operating income would increase by 58,500(11700+46800).It is because the cost of goods sold would decrease by the amount allocated to work in process (11700) and finished goods (46800).

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY HITTING ??

THANK YOU